OpenStudy (anonymous):

Your station charges $16.50 for a lubrication job. As a promotion you sell six coupons for lubrication jobs for $90. What percentage discount are you offering for customers who purchase the 6-coupon lube book?

OpenStudy (anonymous):

90/6*16.5 =90/99 =.909 ans*100 =you are paying 90.9% of full price which would then mean you would be saving 10.1%

OpenStudy (triciaal):

another approach 6 jobs @ 16.50 each would cost 99 6 jobs with coupon cost 90 discount (99 - 90) = 9 percent means per 100 discount as a percentage ,x 9/99 = x/100 x = 900/99 = 9%